OpenStudy (anonymous):

20=500/(10+n) solve for n.

OpenStudy (anonymous):

start with \[20(10+n)=500\] and solve that one

OpenStudy (anonymous):

20(10+n)=500 200-200+n=500-200?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[20(10+n)=500\]\[200+20n=500\] \[20n=300\]\[n=300\div 20=30\div 2=15\]

OpenStudy (anonymous):

Thank You now i understand!